Here is a standard tire gauge, shown as a cross-se
Answer #1

spring constant k = 9.24 x 102 N/m

when the rod moves out through a distance 2.0 cm, compression in the spring x is 2.0 cm.

Tension in the spring is T = kx = 18.48 N

Area of piston A = 7.9 x 10-5 m2

(a) differential pressure Pd = T/A = 2.34 x 105 N/m2 = 2.34 x 105 Pa = 33.9 psi

(b) actual pressure = differential pressure + atmospheric pressure = 33.9 + 14.7 = 48.6 psi